OpenStudy (anonymous):

i could really use some help factoring radical expressions. taking my first math class since high school, a little behind.

OpenStudy (anonymous):

you got one in mind?

OpenStudy (anonymous):

well, why does the square root of 6x*thee square root of 3x become 3x*the square root of 2

OpenStudy (anonymous):

\[6x\sqrt{3x}\]?

OpenStudy (anonymous):

it doesn't i must have misread the question

OpenStudy (anonymous):

try writing like sqrt(2x) for example, then it will be easier to read

OpenStudy (anonymous):

oh maybe \[\sqrt{6x}\times \sqrt{3x}\] like that?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

ok it goes like this \[\sqrt{6x}\times \sqrt{3x}=\sqrt{6x\times 3x}=\sqrt{18x^2}=\sqrt{9\times x^2\times 2}\] \[=\sqrt{9}\times \sqrt{x^2}\times \sqrt{2}=3x\sqrt{2}\] a little more drawn out then necessary but i did not want to skip any steps

OpenStudy (anonymous):

thanks for the reply. do you know any good sights for video lectures?